Please solve for Growth rate of real GDP and show work

Given below are data on real GDP and potential GDP for the nation of Anaziland for the years 2009–2013, in billions of 2009 currency. For each year, calculate the output gap as a percentage of potential GDP and state whether the gap is a recessionary gap or an expansionary gap. Also calculate the year-to-year growth rates of real GDP.

Instructions: Enter your response as a percentage rounded two decimal places.If you are entering any negative numbers be sure to include a negative sign (-) in front of those numbers.

Year

Real GDP

Potential GDP

Output gap

Type of gap

Growth rate of real GDP

2009

13,250

13,660

-3.00%

recessionary

-----

2010

13,990

13,300

5.19%

expansionary

%

2011

14,450

13,400

7.84%

expansionary

%

2012 13,880 14,060 -1.28% recessionary %
2013

14,080

14,500

-2.90% recessionary %
Answer #1

Growth rate of real GDP

2010 = ((13,990-13.250)/13,250))*100= (740/13,250)x100=5.58%

2011 = ((14,450-13,990)/13,990)*100= (460/13,990)x100=3.29%

2012=((13,880-14,450)/14,450*100= (-570/14,450)x100=-3.94%

2013=((14,080-13,880)/13,880*100= (200/13,880)x100=1.44%%